OpenStudy (anonymous):

it says graph 7x - 4y = 28, using the intercept method

OpenStudy (anonymous):

i'm not sure how to break this down

OpenStudy (hoblos):

so you have to find the x-intercept and the y-intercept to find x-intercept replace y=0 to find y-intercept replace x=0 can you do that by yourself??

OpenStudy (anonymous):

so 7(0)=0 and 4(0)=0?

OpenStudy (anonymous):

this stuff is so confusing to me

OpenStudy (hoblos):

for y=0 7x = 28 x=4 then x-intercept is (4,0) for x=0 -4y=28 y=-7 then y-intercept is (0,-7) now plot these two points and join them

OpenStudy (hoblos):

do you get it now??

OpenStudy (anonymous):

this is going to take some time

OpenStudy (hoblos):

why?? do you know how to plot the two points that i gave you

OpenStudy (anonymous):

i'm trying to figure it out

OpenStudy (anonymous):

ok i think i got that one right

OpenStudy (anonymous):

Ok, now I have a y = 4/5x + 4. How would I graph that one

OpenStudy (hoblos):

if you replace x=0 in the equation then y=?? try it and tell me what you get

OpenStudy (anonymous):

so 5x would turn in to 0?

OpenStudy (hoblos):

mm wait.. is it 4 4 ------- or -------- x 5x 5

OpenStudy (anonymous):

y = 4/5x + 4

OpenStudy (hoblos):

you mean 4 ------ 5x +4

OpenStudy (hoblos):

soo yes 5x=0 what is y then?

OpenStudy (anonymous):

0?

OpenStudy (hoblos):

y= 4/(0+4) = 4/4 = 1

OpenStudy (anonymous):

i wouldnt even know how to graph that

OpenStudy (hoblos):

are you sure that the equation you have is of the form 4 ------ 5x +4

OpenStudy (anonymous):

y = 4/5x + 4

hero (hero):

Vince, you find the x and y intercepts because they are two points that you can plot on an xy plane. Since between any two points is a straight line, you can simply draw a line through those points and label it y = 4x/5 + 4

OpenStudy (hoblos):

can you write it using Draw so that it become clear